OpenStudy (anonymous):

y=(m^2)x+(1/m) is the tangent to the parabola y^2 = 16x for what value of m?

OpenStudy (anonymous):

sub first in the second, expand out, you gert quadratic in x , find discriminant and set to zero

OpenStudy (anonymous):

putting the equation of tangent in the eqn. of parabola we get (m^2x+1/m)^2=16x m^4x^2(2m-16)x+1/m^2=0. as it is a tangent so discriminant is 0. (2m-16)^2-4m^2=0 gives you m=4.

OpenStudy (anonymous):

please check it.....

OpenStudy (anonymous):

the way you type is very annoying

OpenStudy (anonymous):

should put brackets around stuff

OpenStudy (anonymous):

all right. thanx you guys, that helped.

OpenStudy (anonymous):

elecengineer .....i did it to make him understand the problem.....